Hi there!
The angle of elevation from me to the top of a library is 15* (degrees). We know that the library is 50 feet tall.
a) How far am I standing from the library?
b) How long of a rope would I need to stretch from me to the top of the library?

Answer by Boreal(15235)

|
|50
|--------------------------15 deg
|=======================You
---------------y-----------------
tan 15=50/y
y=50/tan 15=186.6 feet distant
The rope would be the hypotenuse. One could use the pythagorean theorem, but sin 15=50/x
x=50/sin 15=193.2 feet
check: 50^2+186.6^2=2500+34819.56=37319.56
sqrt (37319.56)=193.2